OpenStudy (anonymous):

Explain, in complete sentences, how you would factor 32x6 – 162.

jhonyy9 (jhonyy9):

2*16x6 - 2*81=2(16x6 -81)=2(4x3 -9)(4x3 +9)

jhonyy9 (jhonyy9):

\[32x ^{6} -162=2(16(x ^{3})^{2} -81)=2((4x ^{3})^{2} -9^{2})=\]

jhonyy9 (jhonyy9):

\[=2(4x ^{3} -9)(4x ^{3} +9)\]